# Knowledge Distillation for Underwater Feature Extraction and Matching via GAN-synthesized Images

### v1

@@ Abstract
Autonomous Underwater Vehicles (AUVs) play a crucial role in underwater exploration. Vision-based methods offer cost-effective solutions for localization and mapping in the absence of conventional sensors like GPS and LIDAR. However, underwater environments present significant challenges for feature extraction and matching due to image blurring and noise caused by attenuation, scattering, and the interference of marine snow. In this paper, we aim to improve the robustness of the feature extraction and matching in the turbid underwater environment using the cross-modal knowledge distillation method that transfers the in-air feature extraction models to underwater settings using synthetic underwater images as the medium. We first propose a novel adaptive GAN-synthesis method to estimate water parameters and underwater noise distribution, to generate environment-specific synthetic underwater images. We then introduce a general knowledge distillation framework compatible with different teacher models. The evaluation of GAN-based synthesis highlights the significance of the new components, i.e. GAN-synthesized noise and forward scattering, in the proposed model. Additionally, the downstream application of feature extraction and matching (VSLAM) on real underwater sequences validates the effectiveness of the transferred model.
@@ I INTRODUCTION
Recently, Autonomous Underwater Vehicles (AUVs) have demonstrated great potential in underwater explorations, replacing human divers in hazardous and demanding scenarios. They are capable of performing various tasks, such as underwater facility repairs and inspections. Accurate localization and mapping are imperative for the successful execution of these missions, presenting a significant challenge due to the unavailability of in-air sensors like the Global Positioning System (GPS) and LIDAR in underwater environments. Consequently, AUVs have to depend on alternative sensors, such as cameras and sonar, for environmental perception and position estimation. Vision-based methods, which utilize cameras as sensing equipment, stand out for their affordability and ease of implementation [1].
Feature point extraction and matching is essential for deriving 3D geometric information from 2D images in applications like Structure-from-Motion (SfM) and Visual Simultaneous Localization and Mapping (VSLAM). Classical methods, such as SIFT [2] and ORB [3], are widely used; for instance, COLMAP [4] employs SIFT to identify correspondences between image frames for map reconstruction. Similarly, the ORB algorithm [3] is the front end of the ORB-SLAM [5] framework, facilitating camera motion estimation across frames. Recently, learning-based methods have demonstrated significant potential in feature extraction and description [6, 7, 8, 9], particularly in challenging scenarios such as fast motion [10], underwater turbidity [11], and low-light environments [12].
@@
To enhance feature extraction and matching in blurry underwater images, several studies have implemented image restoration techniques to improve image quality and enhance the finding of feature correspondences [17, 18]. In [19], an additional classification model is used to mitigate incorrect feature detections on marine snow. Another promising approach involves cross-modal knowledge distillation. Earlier work, UFEN [11], leverages synthetic underwater images as a medium to transfer in-air feature extraction and matching knowledge to underwater scenarios. However, in UFEN, the synthetic images are generated using randomly sampled water parameters from various water types [20], with image noise simplified as white noise, limiting its adaptability to diverse underwater environments with varying characteristics.
@@
The Generative Adversarial Network (GAN) [21] has been widely used in underwater image generation tasks to enhance the similarity between synthesized images and target underwater scenes, addressing the challenges of groundtruth collection in underwater environments. This approach serves as a precursor for various downstream tasks, including image restoration [22, 23, 24, 25] and depth estimation [26, 27]. However, synthetic images for feature extraction should account for not only color appearance in underwater environments but also critical factors such as blurring and image noise, which greatly impact the robustness of feature extraction and matching in underwater settings.
To address these challenges, we propose an adaptive GAN-based synthesis method integrated with a generalized knowledge distillation framework, enabling in-air feature extraction and description models to adapt to underwater scenarios. First, we propose a GAN-based approach that incorporates adaptive image noise distribution and forward scattering into an image formation model to generate synthetic underwater images for the target environment (see Section III). Building upon previous work, we then present a framework that allows one to choose different in-air feature extraction networks as teacher models by introducing adaptation steps for different descriptor formats and dimensions (see Section IV). The overall framework is illustrated in Fig. 1. The key contributions of this paper are as follows:
@@
  • We validate the effectiveness of our approach by integrating the transferred model into a VSLAM framework, improving robustness in challenging underwater environments.
@@ II-A Underwater Imaging Formation Model
Underwater imaging differs greatly from in-air imaging. Recent studies have streamlined the underwater imaging model into two main components: the direct signal and the backscattering signal [28, 29, 30, 31]. The basic model can be expressed as:
@@
here, $I_{c}(x)$ and $J_{c}(x)$ denote the intensities at pixel $x$ in the captured underwater image and the unattenuated image, respectively, across the $c\in\{R,G,B\}$ color channels. The direct signal describes photons traveling directly from objects to the camera, experiencing energy loss due to scattering and absorption, with the attenuation rate determined by the beam attenuation coefficient $\beta_{c}$ and the distance to the camera, $z(x)$. In contrast, backscattering is caused by the reflection of ambient light $B_{c}^{\infty}$ from particles in water.
However, the basic model in (1) is overly simplified, overlooking characteristics of underwater imaging, such as forward scattering and the phenomenon known as marine snow. Marine snow, composed of organic material, detritus, and inorganic particles, plays a crucial role in underwater imaging. Its presence can lead to incorrect feature detections in underwater images [19] and inconsistent feature descriptions [11]. Modeling and synthesizing marine snow is particularly challenging due to its highly variable shapes and structures. Some studies address this challenge by employing mathematical models to simulate its behaviour [32, 33]. Furthermore, the motion blur caused by the movement of marine snow adds additional complexity to the synthesis.
This paper adopts the model in (1) as the foundation for synthetic image generation. The refined version of this model is introduced in Section III.
II-B Learning-based Feature Extraction and Description
In recent years, learning-based feature extraction and matching networks have demonstrated promising performance in identifying correspondences between image frames. Score map-based methods, which generate a score map to identify potential feature points and a descriptor map to assign vector descriptions to these points, represent one main category [34]. SuperPoint, proposed in [6], is trained starting from corner points on synthetic shapes and refined through a homographic adaptation strategy. R2D2 [9] focuses on discriminative learning the repeatability and reliability of features to achieve reliable detection and description. ALIKE [34] and ALIKED [8] utilize a differentiable keypoint detection (DKD) and dispersity peak loss for accurate and repeatable keypoint. Meanwhile, learning-based feature extraction methods have been employed to enhance the robustness of feature matching in challenging environments [10, 11]. The Underwater Feature Extraction Network (UFEN) [11] enhances feature extraction and matching from blurry underwater images by generating synthetic images and applying cross-modal knowledge distillation
@@ II-C Underwater Image Synthesis
Acquiring underwater ground truth is challenging for tasks like color restoration and depth estimation. To address this, synthetic underwater images are often used to build large underwater datasets [22, 26, 23]. The foundational approach leverages the underwater imaging formation model, using in-air RGB-D images with known water-type parameters to generate synthetic underwater images [11, 35]. The GAN method is an approach capable of synthesizing the target underwater environment. WaterGAN, proposed in [22], takes RGB-D images as input to generate synthetic underwater images by combining GAN methods with the underwater imaging formation model. These synthetic images serve as ground truth for image restoration tasks. Recent work [27] employs a diffusion model to generate diverse underwater scenes using in-air depth images as input, targeting underwater monocular depth estimation tasks.
@@ III GAN-Based Image Synthesis
In this section, we propose a model-based GAN approach used for underwater image synthesis, including the refined imaging formation model we employed and the GAN framework for estimating water parameters and noise distribution.
@@ III-A1 Forward Scattering
In the model (1), the forward scattering component, which describes light scattering at small angles biased toward the line of sight, is omitted, as its influence on image degradation is less than that of the other two components [36]. However, the existence of forward scattering causes blurriness in images, affecting feature detection and description. This scattering effect can be represented by the Point Spread Function (PSF) applied to the direct signal [37, 36]. Therefore we consider the following image model with forward scattering in this paper:
@@ III-A2 Image Noise
Image noise from marine snow in underwater environments affects feature extraction. A recent study [19] demonstrated that marine snow significantly impairs feature extraction performance. Moreover, the irregular motion of marine snow leads to inconsistent feature descriptions across image frames, thereby reducing feature-matching accuracy. Our synthesis process thus accounts for the image noise caused by marine snow, whose effect accumulates with increasing scene distance, similar to backscattering. We define the noise map $M$, which has the same dimensions as the image, applied to the pure backscattering background $B_{c}^{\infty}$. The backscattering background can be viewed as being partially filled with marine snow, represented as $B_{c}^{\infty}\bigl{(}1-M(x)\bigr{)}+M(x)$. Here, $M(x)\in[0,1]$ represents a weight value at each pixel location $x$. We denote the additional image noise as $B_{c}^{Noise}$, formulated by:
@@ III-B2 GAN Image Generation
Figure 3: Training the Noise Generator $\mathbf{G}$: The training of the generator uses a discriminator $\mathbf{D_{d}}$ to ensure a uniform distribution of the generated noise map.
@@ III-B3 Evenly Distributed Noise Constraint
Meanwhile, to ensure that the generated noise map $M$ from $\mathbf{G}$ is evenly distributed across images, we introduce an additional discriminator $\mathbf{D_{d}}:\mathbb{R}^{m\times w\times h}\to[0,1]$, in which $m$ is the batch size, $w$ and $h$ are the grid size in width and height. Let $\mathcal{G}$ denote the set of grid locations, uniformly sampled from the noise map $M$. For each batch of noise samples =$\mathcal{M}=\{M_{1},M_{2},\dots,M_{m}\}$, with each $M_{j}\sim\mathbb{M}$ $(j=1,\dots,m)$, the discriminator compares patches sampled from the same grid location across the batch with patches sampled from random grid locations across all batch samples. Let $g_{i}\in\mathcal{G}$ be the grid location on patch $i$ ($i=1,\dots,\frac{HW}{hw}$), $W$ and $H$ are the image size in width and height. The batch of patches from the same grid location is denoted as:
@@ IV Adaptive knowledge distillation framework for underwater feature extraction
After generating synthetic underwater images, we perform knowledge distillation to transfer a pre-trained in-air feature extraction model to the underwater domain.
@@ IV-B Dispersity peak loss
To improve score precision at detected feature points, we apply the dispersity peak loss [34]. Given an image $I_{c}^{Noise}$, $N^{Stu}$ feature points $x_{i}^{Stu}$ ($i\in\{1,\dots,N^{Stu}\}$) are extracted using $\mathcal{X}^{Stu}$. For each feature point $\mathbf{x}^{Stu}_{i}$, a local patch of size $S\times S$ centered at $\mathbf{x}^{Stu}_{i}$ is defined and the Softmax [34] is applied to the score map within these patches. The pixel locations in the patch centered at $x_{i}^{Stu}$ are denoted as $x_{i,j}^{Stu}$, where $j\in\{1,2,\dots,S^{2}\}$, and the scores after Softmax at these pixels are denoted as $s(x_{i}^{Stu})$. $r(x_{i}^{Stu},x_{i,j}^{Stu})$ is the distance function, same to that in (4). Dispersity peak loss $\mathcal{L}_{peak}$ is defined as:
@@ IV-C Matching loss
where $Z$ is a scaling factor, equal to 1 for the float descriptor or to the descriptor dimension in the binary case. In (16), $\text{dist}(\cdot)$ represents the descriptor distance between the matching descriptors $d_{i}^{Stu}$ and the corresponding descriptor in the set $d_{j}^{\mathcal{H}}$, while $\text{dist}_{n}(\cdot)$ represents the minimal descriptor distance (e.g., $L_{2}$ distance) to the non-matching descriptors in $d_{j}^{\mathcal{H}}$. $P$ and $Q$ are two distance margins for the matching and non-matching descriptors, respectively. To accommodate the varying descriptor dimensions and formats across teacher models, we calculate the binary descriptor distance using the following equation:
@@
Meanwhile, same as the binarization step in UFEN [11], to prevent zero gradients during backpropagation in training, we employ the straight-through estimator.
@@ IV-D Final Loss
We use the weight parameters $\gamma_{1}$ and $\gamma_{2}$ to adjust the weights of the different training losses.
@@ V Experiments
(a) Pure background 1
(b) Pure background 2
Figure 4: Naive GAN Testing on Pure Background Images: In the two displayed subsets, the left image represents the captured real image, while the right image is synthetic.
@@ V-A Implementation Details
The training dataset for the GAN synthesis process consists of 1449 images from the NYU-v2 dataset [41], with 1149 images used for training and 300 for testing. The depth values are normalized between 0.5m and 5m to ensure scene consistency. We first perform the parameter estimation of the formation model in (2). The training is conducted with a batch size of 4 for 10k iterations. The learning rate is set to $1\times 10^{-3}$ for the trainable parameters and $1\times 10^{-4}$ for the discriminator $\mathbf{D}$, with updates occurring every 5 iterations. Then, the noise distribution estimation is performed with a batch size of 4 for 20k iterations. The learning rate is set to $1\times 10^{-5}$ for both the generator $\mathbf{G}$ and the reinitialized discriminators $\mathbf{D}$ and $\mathbf{D_{d}}$, with updates every 10 iterations, and $\lambda$ is set to 0.1. The structure of the discriminators follows the multi-scale discriminators proposed in [42]. The generator, which is based on [43], replaces the encoder with a fully connected layer to adjust input random noise with size $N=5$ to the input size of the residual blocks and modifies the output layer to employ a Sigmoid activation function. In the knowledge distillation step, we continue using the dataset from the previous stage, training for 30 epochs. The learning rate is set to $1\times 10^{-5}$ for ALIKE and $1\times 10^{-6}$ for SuperPoint. $\gamma_{1}$, $\gamma_{2}$, and $\alpha_{KD}$ are set to 1, 1, and 0.01, respectively.
(a) $J_{c}$
(b) $I_{c}$
(c) $I^{Nosie}_{c}$
(d) $T_{c}$

### v2

@@ Abstract
Autonomous Underwater Vehicles (AUVs) play a crucial role in underwater exploration. Vision-based methods offer cost-effective solutions for localization and mapping in the absence of conventional sensors like GPS and LiDAR. However, underwater environments present significant challenges for feature extraction and matching due to image blurring and noise caused by attenuation, scattering, and the interference of marine snow. In this paper, we aim to improve the robustness of the feature extraction and matching in the turbid underwater environment using the cross-modal knowledge distillation method that transfers the in-air feature extraction and matching models to underwater settings using synthetic underwater images as the medium. We first propose a novel adaptive GAN-synthesis method to estimate water parameters and underwater noise distribution, to generate environment-specific synthetic underwater images. We then introduce a general knowledge distillation framework compatible with different teacher models. The evaluation of GAN-based synthesis highlights the significance of the new components, i.e. GAN-synthesized noise and forward scattering, in the proposed model. Additionally, VSLAM, as a representative downstream application of feature extraction and matching, is employed on real underwater sequences to validate the effectiveness of the transferred model. Project page: https://github.com/Jinghe-mel/UFEN-GAN.
Index Terms:
Underwater image synthesis, knowledge distillation, feature extraction and matching, underwater VSLAM
@@ I INTRODUCTION
Recently, Autonomous Underwater Vehicles (AUVs) have demonstrated great potential in underwater explorations, replacing human divers in hazardous scenarios. They are capable of performing various tasks, such as underwater facility repairs and inspections. Accurate localization and mapping are imperative for the successful execution of these missions, presenting a significant challenge due to the unavailability of in-air sensors like the Global Positioning System (GPS) and LiDAR. Consequently, AUVs depend on alternative sensors, such as cameras and sonar, for environmental perception and position estimation. Vision-based methods, which utilize cameras as sensing equipment, stand out for their affordability and ease of implementation [1].
Feature point extraction and matching is essential for deriving 3D geometric information from 2D images in applications like Structure-from-Motion (SfM) and Visual Simultaneous Localization and Mapping (VSLAM). Classical methods, such as SIFT [2] and ORB [3], are widely used; for instance, COLMAP [4] employs SIFT to identify correspondences between image frames for map reconstruction. Similarly, ORB [3] is the front end of the ORB-SLAM [5] framework, facilitating camera motion estimation across frames. Recently, learning-based methods have demonstrated significant potential in feature extraction and description [6, 7, 8, 9], particularly in challenging scenarios such as fast motion [10], underwater turbidity [11], and low-light environments [12].
@@
To enhance feature extraction and matching in blurry underwater images, several studies use image restoration techniques to improve image quality and enhance the finding of feature correspondences [17, 18]. In [19], an additional classification model is used to mitigate incorrect feature detections on marine snow, which is composed of organic material, detritus, and inorganic particles in water. Another promising approach involves cross-modal knowledge distillation [20], which transfers feature extraction and matching knowledge learned on clear in-air images to models operating on visually distinct underwater domains. For example, UFEN [11] employs synthetic underwater images as a medium for this knowledge transfer.
@@
The Generative Adversarial Network (GAN) [21] has been widely used in underwater image generation tasks to enhance the similarity between synthesized images and target underwater scenes, addressing the challenges of groundtruth collection in underwater environments. This approach serves as a precursor for various downstream tasks, including image restoration [22, 23, 24, 25] and depth estimation [26, 27]. However, synthetic images for feature extraction should account for not only color appearance in underwater environments but also critical factors such as blurring and image noise, which greatly impact the robustness of feature extraction and matching in underwater settings. While the GAN-based enhancement method proposed in [28] improves degraded underwater images and facilitates VSLAM, it may also unintentionally amplify the visibility of floating marine snow, which may lead to a degradation in feature matching performance.
To address these challenges, we propose an adaptive GAN-based synthesis method integrated with a generalized knowledge distillation framework, enabling in-air feature extraction and description models to adapt to underwater scenarios. First, unlike UFEN, which generates the synthetic images using randomly sampled water parameters from various water types [29], and injects visually unrealistic Gaussian noise into synthetic images, we propose a GAN-based approach that incorporates adaptive noise distribution and forward scattering into the image formation. We further demonstrate the necessity and effectiveness of these components for improving feature matching transfer performance (see Section III). Building upon previous work, UFEN, which is specifically designed for transferring SuperPoint [6], we propose a general cross-modal knowledge distillation framework that can adapt to various score map-based feature extraction and matching models, supporting different output structures and descriptor formats (see Section IV).
The overall framework is illustrated in Fig. 1. The key contributions of this paper are as follows:
@@
  • Our experiments demonstrate that incorporating forward scattering and GAN-synthesized noise significantly improves the transfer performance of feature matching in underwater environments.
  • We validate the effectiveness of our method by integrating the transferred model into the VSLAM framework.
@@ II-A Underwater Imaging Formation Model
Recent studies have streamlined the underwater imaging model into two components: the direct and the backscattering signal [30, 31, 32, 33], which can be expressed as:
@@
where $I_{c}(x)$ and $J_{c}(x)$ denote the intensities at pixel $x$ in the captured underwater image and the unattenuated image, respectively, across the $c\in\{R,G,B\}$ color channels. The direct signal describes photons travelling directly from objects to the camera, experiencing energy loss due to scattering and absorption, with the attenuation rate determined by the beam attenuation coefficient $\beta_{c}$ and the distance to the camera, $z(x)$. In contrast, backscattering is caused by the reflection of ambient light $B_{c}^{\infty}$ from particles in water. However, the basic model in (1) is overly simplified, overlooking characteristics of underwater imaging, such as forward scattering and the phenomenon known as marine snow. Marine snow, composed of organic material, detritus, and inorganic particles, plays a crucial role in underwater imaging. Its presence can lead to incorrect feature detections in underwater images [19] and inconsistent feature descriptions [11]. Modelling and synthesizing marine snow is particularly challenging due to its highly variable shapes and structures. Some studies address this challenge by employing mathematical models to simulate its behaviour [34, 35]. Furthermore, the motion blur caused by the movement of marine snow adds additional complexity to the synthesis. This paper adopts the model in (1) as the foundation for synthetic image generation.
II-B Learning-based Feature Extraction and Matching
In recent years, learning-based feature extraction and matching networks have demonstrated promising performance in identifying correspondences between image frames. Score map-based methods, which generate a score map to identify potential feature points and a descriptor map to assign descriptions to these points, represent one main category. SuperPoint, proposed in [6], is trained starting from corner points on synthetic shapes and refined through a homographic adaptation strategy. R2D2 [9] jointly learns the repeatability and reliability of features using a discriminative loss, enabling reliable detection and description. ALIKE [36] and ALIKED [8] utilize a differentiable keypoint detection (DKD) and dispersity peak loss for accurate and repeatable keypoint detection. Meanwhile, learning-based feature extraction methods have been employed to improve feature matching robustness in challenging environments [10, 11], such as UFEN [11], which addresses underwater blur by generating synthetic images and applying cross-modal knowledge distillation.
@@ II-C Underwater Image Synthesis
Acquiring underwater groundtruth is challenging for tasks like color restoration and depth estimation. To address this, synthetic underwater images are often used to build large underwater datasets [22, 26, 23]. The foundational approach leverages the underwater imaging formation model, using in-air RGB-D images with known water-type parameters to generate [11, 37]. The GAN method is an approach capable of synthesizing the target underwater environment. WaterGAN [22] takes RGB-D images as input to generate synthetic images by combining GAN with the imaging formation model. These synthetic images serve as groundtruth for image restoration tasks. Recent work [27] employs a diffusion model to generate diverse underwater scenes using depth images as input, targeting underwater monocular depth estimation tasks.
@@ III GAN-Based Image Synthesis
In this section, we propose a model-based GAN approach used for underwater image synthesis, including the refined imaging formation model and the GAN framework.
@@ III-A1 Forward Scattering
In the model (1), the forward scattering component, which describes light scattering at small angles biased toward the line of sight, is omitted, as its influence on image degradation is less than that of the other two components [38]. However, the existence of forward scattering causes blurriness in images, affecting feature detection and description. This scattering effect can be represented by the Point Spread Function (PSF) applied to the direct signal [39, 38]. Therefore, we consider the following image model with forward scattering in this paper:
@@ III-A2 Image Noise
Image noise from marine snow affects feature extraction. A recent study [19] demonstrated that marine snow significantly impairs feature extraction and matching performance. Moreover, the irregular motion of marine snow leads to inconsistent feature descriptions, thereby reducing feature-matching accuracy. Our synthesis process thus accounts for the image noise caused by marine snow, which consists of suspended particles that are part of backscattering and accumulate with scene depth. We model this noise as an additive perturbation to the pure backscattering background, similar to [19]. We define the noise map $M$, which has the same dimensions as the image, applied to the pure backscattering background $B_{c}^{\infty}$. The backscattering background can be viewed as being partially filled with marine snow, represented as $B_{c}^{\infty}\bigl{(}1-M(x)\bigr{)}+M(x)$. Here, $M(x)\in[0,1]$ represents a weight value at each pixel location $x$. We denote the additional image noise as $B_{c}^{Noise}$, formulated by:
@@ III-B2 GAN Image Generation
Figure 3: Training the Noise Generator $\mathbf{G}$: The discriminator $\mathbf{D_{d}}$ is to ensure a uniform distribution of the noise map.
@@ III-B3 Evenly Distributed Noise Constraint
Meanwhile, to ensure that the generated noise map $M$ from $\mathbf{G}$ is evenly distributed across images, we introduce an additional discriminator $\mathbf{D_{d}}:\mathbb{R}^{m\times w\times h}\to[0,1]$, in which $m$ is the batch size, $w$ and $h$ are the grid size in width and height. Let $\mathcal{G}$ denote the set of grid locations, uniformly sampled from the noise map $M$. For each batch of noise samples $\mathcal{M}=\{M_{1},M_{2},\dots,M_{m}\}$, with each $M_{j}\sim\mathbb{M}$ $(j=1,\dots,m)$, the discriminator compares patches sampled from the same grid location across the batch with patches sampled from random grid locations across all batch samples. Let $g_{i}\in\mathcal{G}$ be the grid location on patch $i$ ($i=1,\dots,\frac{HW}{hw}$), $W$ and $H$ are the image size in width and height. The batch of patches from the same grid location is denoted as:
@@ IV Adaptive knowledge distillation framework for underwater feature extraction
After generating synthetic underwater images, we apply cross-modal knowledge distillation, where the supervision for feature matching is provided by the teacher model on in-air images to train the student model on the corresponding synthetic underwater images. Specifically, the teacher model takes in-air images as input, whereas the student model takes the corresponding synthetic images as input.
@@ IV-B Dispersity Peak Loss
To improve score precision at detected feature points, we apply the dispersity peak loss [36]. Given an image $I_{c}^{Noise}$, $N^{Stu}$ feature points $x_{i}^{Stu}$ ($i\in\{1,\dots,N^{Stu}\}$) are extracted using $\mathcal{X}^{Stu}$. For each feature point $\mathbf{x}^{Stu}_{i}$, a local patch of size $S\times S$ centered at $\mathbf{x}^{Stu}_{i}$ is defined and the Softmax is applied to the score map within these patches. The pixel locations in the patch centered at $x_{i}^{Stu}$ are denoted as $x_{i,j}^{Stu}$, where $j\in\{1,2,\dots,S^{2}\}$, and the scores after Softmax at these pixels are denoted as $s(x_{i}^{Stu})$. $r(x_{i}^{Stu},x_{i,j}^{Stu})$ is the distance function, same to that in (4). Dispersity peak loss $\mathcal{L}_{peak}$ is defined as:
@@ IV-C Matching Loss
where $Z$ is a scaling factor, equal to 1 for the float descriptor or to the descriptor dimension in the binary case. In (16), $\text{dist}(\cdot)$ represents the descriptor distance between the matching descriptors $d_{i}^{Stu}$ and the corresponding descriptor in the set $d_{j}^{\mathcal{H}}$, while $\text{dist}_{n}(\cdot)$ represents the minimal descriptor distance (e.g., $L_{2}$ distance) to the non-matching descriptors in $d_{j}^{\mathcal{H}}$. $P$ and $Q$ are two distance margins for the matching and non-matching descriptors, respectively. To accommodate the varying descriptor dimensions and formats across teacher models, we calculate the binary descriptor distance using:
@@
Following UFEN [11], we use the straight-through estimator to avoid zero gradients during backpropagation in binarization.
@@ IV-D Final Loss
The weights $\gamma_{1}$ and $\gamma_{2}$ are used to balance the loss terms.
@@ V-A Implementation Details
The training dataset for the GAN synthesis process consists of 1449 images from the NYU-v2 dataset [43], with 1149 images used for training and 300 for testing. The depth values are normalized between 0.5m and 5m to ensure scene consistency. Inspired by the adversarial disentanglement training method in [44], we first perform the parameter estimation of the formation model in (2) by freezing the noise generator and optimizing only the physical parameters in the formation model. The training is conducted with a batch size of 4 for 10k iterations. The learning rate is set to $10^{-3}$ for the trainable parameters and $10^{-4}$ for the discriminator $\mathbf{D}$, with updates occurring every 5 iterations. Then, the noise distribution estimation is performed with a batch size of 4 for 20k iterations. The learning rate is set to $10^{-5}$ for both the generator $\mathbf{G}$ and the reinitialized discriminators $\mathbf{D}$ and $\mathbf{D_{d}}$, with updates every 10 iterations, and $\lambda$ is set to 0.1. The structure of the discriminators follows the multi-scale discriminators proposed in [45]. The generator, which is based on [46], replaces the encoder with a fully connected layer to adjust input random noise with size $N=10$ to the input size of the residual blocks and modifies the output layer to employ a Sigmoid activation function. In the knowledge distillation step, we continue using the dataset from the previous stage, training for 30 epochs. The learning rate is set to $10^{-5}$ for ALIKE and $10^{-6}$ for SuperPoint. $\gamma_{1}$, $\gamma_{2}$, and $\alpha_{KD}$ are set to 1, 1, and 0.01, respectively.
TABLE I: Quantitative Evaluation on the Lake-C and EASI-T Sequences
Models Lake-C EASI-T AQUALOC
M. N M. R M. N M. R M. N M. R
ALIKE 258.0 0.877 139.5 0.641 212.2 0.756
ALIKE (KD) 264.0 0.929 177.6 0.769 214.0 0.817
SuperPoint 153.3 0.592 130.1 0.550 306.5 0.904
SP (KD) 275.2 0.904 191.3 0.780 335.4 0.952
UFEN 244.2 0.807 146.0 0.576 300.1 0.890
SP (KD-B) 246.6 0.861 172.9 0.744 333.2 0.943
$J_{c}$
$I_{c}$
$I_{c}^{\text{Noise}}$
$T_{c}$ (real)